Grade 12Physical Chemistry

I am having a problem in the following question.

Ammonium carbamate dissociates as : In a closed vessel containing ammonium carbamate in equilibrium with ammonia and carbon dioxide, ammonia is added such that partial pressure of NH3 now equals to the original total pressure. The ratio of total pressure to the original pressure now is ?

Nh2coonh4(s)----2nh3(g)+co2(g)

I tried solving the problem and I know I am quite close to the answer which is 31/27. Firstly I took out ko for equilibrium assuming total pressure to be p. So kc= (2p/3)^2*p/3=4p^3/27

After this the par pressure of ammonia becomes p so let's say that of co2 is p/3+x. Now since kc doesn't change so p/3+x=4p^2/27(cancelling one of the p) But solving further and finding new pressure and dividing by the old one doesn't give the ans. please help me.

Jai Mahajan
11 Years agoGrade 12

2 Answers

Sunil Kumar FP
11 Years ago
you are correct upto this point
“I tried solving the problem and I know I am quite close to the answer which is 31/27. Firstly I took out ko for equilibrium assuming total pressure to be p. So kc= (2p/3)^2*p/3=4p^3/27“

Then after adding NH3 the equilibrium pressure of NH3 becomes p which is the original pressure.
since Kp will be same for the above reaction.
Therefore
4p^3/27=p^2*x
where x is the equilibrium partial pressure of CO2.
x=4p/27
Thus the total pressure=4p/27+p=31p/27

The ratio is =31p/27/p=31/27
ankit singh
5 Years ago

  NH2COONH4(s)2NH3(g)+CO2(g)
Initial                                                 2P                  P
KP=(PNH3)2(PCO2)
KP=(2P)2(P)     ......(i)
PT(initial)=3P

               NH2COONH4(s)2NH3(g)+CO2(g)
Final                                                  3P                  P
Kp=(3P)2(P)      ......(ii)
From eq. (i) and (ii)
(2P)2(P)=(3P)2(P)
P=94P
PT(Old)PT(New)=3P3P+P=3P3P+94P=2731.